Assume a bank has to keep a 25% Reserve Ratio. If $100 is deposited at the bank, how much can the bank lend out of that deposit? If the loan is deposited at another bank with a 25% Reserve Ratio, how much will be able to be lent then? If every loan was deposited, how much money would be created (added) to the money supply?
Answer #1

how much can the bank lend out of that deposit

=100*(1-25%)=75

how much will be able to be lent then

=75*(1-25%)

=56.25

If every loan was deposited, how much money would be created (added) to the money supply

=(1/25%)*100

=400

the above is answer..